# How does online shopping affect offline price sensitivity?

### v1

@@ A.1.2 Price and Cost Index Calculations
We construct brand-level price indices for each category by aggregating SKU-level price data from offline transactions (Gordon et al. 2013; Hitsch et al. 2021). First, we construct a store-SKU-month level price matrix from the available transactions data. To begin with, this price matrix could contain both outliers and missing values. Any observed price that is more than four times higher than or less than one-quarter of the median price for each SKU (across all store-month observations) is treated as an outlier and labeled as missing, following Hitsch et al. (2021). This process filters out unreasonable prices before imputing any missing values of store-SKU-month price (when there is no sale or there is an outlier). For these missing prices, we use the highest observed price from the previous three months for the same SKU at the same store to impute the missing value (Gordon et al. 2013). Finally, we exclude SKUs for a particular store and week if we have never observed them being sold in that store previously.
Next, we aggregate SKUs to create brand-level price indices by converting all SKU prices to a per-unit basis for comparable units. All product categories (e.g., dry food and cat sands) are measured by price per kilogram (KG) except for hygienic mats, which are measured by price per piece. Following the approach of Hitsch et al. (2021), the brand price for brand $j$ at store $s$ in month $t$, denoted as $\textrm{BrandPrice}_{jst}$, is calculated as the weighted average price per unit across all SKUs in that brand in the defined market. The weights are based on each SKU’s share of offline sales within the brand at that store over the full 30-month data period.
[…]
We then apply a similar approach to construct a store-SKU-month level cost matrix based on wholesale cost data, and aggregate SKU-level costs to create brand-level cost indices that correspond to each price index. These cost indices serve as the instrument for addressing price endogeneity, as discussed in Section §6. Finally, we use Brazil’s Extended National Consumer Index (IPCA) to obtain inflation-adjusted prices and costs (Brazil Institute of Geography and Statistics (2024) IBGE), with January 2019, the first month of observation, as the base period.
@@ B.1 Supplementary Result: Percentile Price Paid
$\mathcal{I}\{\textrm{Adopter}_{i}\}\times\mathcal{I}\{t\geq T_{i}^{\textrm{adopt}}\}$1.437∗∗ -3.318∗∗ -0.9569 -4.104∗∗∗
@@ E.1 Detailed Results for the Role of Price Difference Across Channels
[…]
@@ E.2.1 Results for Alternate Demand Specification with Binary Adoption Variable
$+\alpha_{3}\mathcal{I}\{\textrm{BuyLastMonth}_{ijt}\}+\beta_{1}\textrm{CF\_Price}_{jt}+\beta_{2}\textrm{CF\_Price\_Post\_OnlineAdoption}_{ijt}+\delta_{ij}~{}~{}\forall j\in\{1,\ldots,J\}$
[…]
@@ E.2.2 Results for Alternate Demand Model Controlling for State Dependence due to Online Adoption
$+\alpha_{3}\mathcal{I}\{\textrm{BuyLastMonth}_{ijt}\}+\alpha_{4}\mathcal{I}\{\textrm{BuyLastMonth}_{ijt}\}\times\textrm{cumPercOnlineSpend}_{it}$
@@ E.3 Detailed results for testing the Parallel Pre-Trend Assumption
A fundamental condition for the credibility of the estimates presented in §$\S$7.3 is the parallel trends assumption. We assess it with an event-study figure following Callaway and Sant’Anna (2021). Figure A7 plots the point estimates and 95% bootstrapped confidence interval for each pre- and post-treatment period, and shows that all pre-treatment coefficients are insignificant from zero. We do not exhibit discernible pre-trends. In addition, the p-values from the Wald tests of the pre-trend assumption are greater than 5% level of significance in each category. We therefore conclude that the parallel trend assumption is reasonably satisfied throughout the analysis.
@@ E.4 Detailed Results for Matching, IPW Approaches to Control for Selection on Observables
Using the predicted propensity score, we apply the R package MatchIt to conduct one-to-one nearest neighbor matching with replacement for each adoption cohort Stuart et al. (2011). For each product category, we first estimate the ATT using the CS based on the matched sample, with results shown in Web Appendix §$\S$§E.4 Table A17. Second, we estimate the ATT using the CS estimator with the inverse propensity score weighting (IPW) method, with results shown in Table A18. These findings are consistent with our main results, demonstrating that consumers become more price sensitive offline after adopting online shopping. Detailed results are below.
@@ E.5 ATT Estimates - Using Not-Yet-Treated Adopters as Control Groups
Appendix F Appendix to Counterfactual Analysis
F.1 Estimation Result for Naive Demand Model
Table A20 provides the estimates for the naive demand model described in 9.1 specified by Equation (20).
Table A20: Coefficient Estimates of the Naive Demand Model
Dry Dog Food Dog Hygiene Cat Hygiene
@@ F.1 Estimation Result for Naive Demand Model
Price -0.3808∗∗∗ (0.0237) -1.5921∗∗∗ (0.1129) -0.2822∗∗∗ (0.0189)
BuyLastMonth 0.1581∗∗∗ (0.0316) -0.1772∗∗∗ (0.0299) 0.3306∗∗∗ (0.0289)
CF_Price 0.1068∗∗∗ (0.0279) 0.5272∗∗ (0.1847) 0.3522∗∗∗ (0.0261)
Customer_Brand_FE Yes Yes Yes
Observations 136,411 78,253 55,410
Log-likelihood -84222.285 -55490.014 -37358.772
@@
F.2 Derivation of Profit Maximizing Uniform Offline Price
First, we derive the market-level profit maximizing uniform offline price for each brand in a given month $t$ that t maximizes category-level offline profit. We drop the index $t$ (month) in the following derivation for simplicity.
Let $\mathcal{M}_{k}$ be a generic demand model, and the superscript $k$ denotes the model used. Let $N$ be the market size, which is the number of offline shoppers who make at least one store visit during the month. The expected profit per customer is:
$R(\mathbf{p};\mathcal{M}_{k})$ $=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{J}(p_{j}-m_{j})\times\frac{\exp(\psi_{ij}(\mathbf{p};\mathcal{M}_{k}))}{1+\sum_{j^{\prime}=1}^{J}\exp(\psi_{ij‘}(\mathbf{p};\mathcal{M}_{k}))}=\frac{1}{N}\sum_{i=1}^{N}R_{i}(\mathbf{p};\mathcal{M}_{k})$ (A12)
where $i$ indexes consumers, $j$ indexes brands, $m_{j}$ is the cost for brand $j$, and $R_{i}(\mathbf{p};\mathcal{M}_{k})$ is customer $i$’s profit contribution given price vector $\mathbf{p}$. Customer $i$’s offline price sensitivity is $\alpha_{i}=\alpha_{i}^{\mathcal{M}_{k}}$ under model $\mathcal{M}_{k}$. For notational convenience, we use $\alpha_{i}$ for the following derivation. Let $e_{ij}=\exp(\psi_{ij}(\mathbf{p};\mathcal{M}_{k}))$ and $\Lambda_{i}(e)=\sum_{j\in[J]}e_{ij}$ under model $\mathcal{M}_{k}$. In our context, $k\in{N,O}$ and we are interested in the uniform pricing under two demand models, including a naive model ($\mathcal{M}_{N}$) defined in Equation (20) and our model ($\mathcal{M}_{O}$) defined in Equation (7).
To find the profit-maximizing price vector $\mathbf{p}_{*}$, we take the first derivative of $R(\mathbf{p};\mathcal{M}_{k})$ with respect to each brand $j$’s price $p_{j}$. The resulting first-order condition (FOC) for the price of brand $j$, $p_{j}$, is:
$\frac{\partial R(\mathbf{p};\mathcal{M}_{k})}{\partial p_{j}}=\frac{1}{N}\sum_{i=1}^{N}\frac{\alpha_{i}e_{ij}}{1+\Lambda_{i}(e)}\left[\frac{1}{\alpha_{i}}+(p_{j}-m_{j})-\frac{\sum_{l=1}^{J}(p_{l}-m_{l})e_{il}}{1+\Lambda_{i}(e)}\right]$ (A13)
The profit-maximizing price vector $\mathbf{p}_{*}$ satisfies the following condition for each brand $j$:
$\frac{1}{N}\sum_{i=1}^{N}\frac{\alpha_{i}e_{ij}}{1+\Lambda_{i}(e)}\left[\frac{1}{\alpha_{i}}+(p_{*,j}-m_{j})-\frac{\sum_{l=1}^{J}(p_{*,l}-m_{l})e_{il}}{1+\Lambda_{i}(e)}\right]=0$ (A14)
Let $Pr_{ij}(\mathbf{p};\mathcal{M}_{k})=\frac{e_{ij}}{1+\Lambda_{i}(e)}.$ be the choice probability of consumer $i$ and brand $j$. Substituting this into the FOC for each brand $j$ yields:
$(p_{*,j}-m_{j})\sum_{i=1}^{N}\alpha_{i}Pr_{ij}(\mathbf{p}_{*};\mathcal{M}_{k})=-\sum_{j=1}^{N}Pr_{ij}(\mathbf{p}_{*};\mathcal{M}_{k})+\sum_{i=1}^{N}\alpha_{i}Pr_{ij}(\mathbf{p}_{*};\mathcal{M}_{k})R_{i}(\mathbf{p}_{*};\mathcal{M}_{k})$ (A15)
$p_{*,j}-m_{j}=-\sum_{i=1}^{N}\frac{Pr_{ij}(\mathbf{p}_{*};\mathcal{M}_{k})}{\sum_{i^{\prime}=1}^{N}\alpha_{i^{\prime}}Pr_{ij^{\prime}}(\mathbf{p}_{*};\mathcal{M}_{k})}+\sum_{i=1}^{N}\frac{\alpha_{i}Pr_{ij}(\mathbf{p}_{*};\mathcal{M}_{k})}{\sum_{i^{\prime}=1}^{N}\alpha_{i^{\prime}}Pr_{ij^{\prime}}(\mathbf{p}_{*};\mathcal{M}_{k})}\times R_{i}(\mathbf{p}_{*};\mathcal{M}_{k})$
$p_{*,j}-m_{j}=c_{j}(\mathbf{p}_{*};\mathcal{M}_{k})+\sum_{i=1}^{N}w_{ij}(\mathbf{p}_{*};\mathcal{M}_{k})R_{i}(\mathbf{p}_{*};\mathcal{M}_{k})$
where $c_{j}(\mathbf{p}_{*};\mathcal{M}_{k})=-\sum_{i=1}^{N}\frac{Pr_{ij}(\mathbf{p}_{*};\mathcal{M}_{k})}{\sum_{i^{\prime}=1}^{N}\alpha_{i^{\prime}}Pr_{ij^{\prime}}(\mathbf{p}_{*};\mathcal{M}_{k})}$ and $w_{ij}(\mathbf{p}_{*};\mathcal{M}_{k})=\frac{\alpha_{i}Pr_{ij}(\mathbf{p}_{*};\mathcal{M}_{k})}{\sum_{i^{\prime}=1}^{N}\alpha_{i}^{\prime}Pr_{ij^{\prime}}(\mathbf{p}_{*};\mathcal{M}_{k})}$. This formulation results in $J$ equations corresponding to the $J$ unknown prices to solve, with the $j_{th}$ row equation given by Equation (22). The optimal price vector is given by $\mathbf{p}_{*}=\mathbf{m}+\mathbf{c}(\mathbf{p}_{*})+\mathbf{W}(\mathbf{p}_{*})^{\prime}\mathbf{R}(\mathbf{p}_{*}).$ We numerically solve for the profit-maximizing price vectors based on a system of $J$ FOC equations.
F.3 Derivation of Profit-maximizing Personalized Offline Prices
We next consider a personalized pricing strategy where the firm can tailor prices to each customer’s elasticity and implement it through targeted coupons and email campaigns. The objective is to maximize each consumer $i$’s expected offline profit, defined as:
$R_{i}(\mathbf{p}^{i};\mathcal{M}_{k})$ $=\sum_{j=1}^{J}(p^{i}_{j}-m_{j})\times\frac{\exp(\psi_{ij}(\mathbf{p}^{i};\mathcal{M}_{k}))}{1+\sum_{j^{\prime}=1}^{J}\exp(\psi_{ij^{\prime}}(\mathbf{p}^{i};\mathcal{M}_{k}))}$ (A16)
where $\mathbf{p}^{i}\in R^{J\times 1}$ represents the personalized price vector for customer $i$. $\psi_{ij}(\mathbf{p}^{i};\mathcal{M}_{k})$ refers to the mean utility of a generic demand model $\mathcal{M}_{k}$. Specifically, $k\in{N,O}$ in our context and we are interested in the personalized pricing under two demand models, including a naive model ($\mathcal{M}_{N}$) defined in Equation (20) and our model ($\mathcal{M}_{O}$) defined in Equation (7).
The first-order condition for maximizing consumer i’s profit with respect to brand $j$’s price is given by:
$p^{i}_{*,j}-m_{j}$ $=-\frac{1}{\alpha_{i}}+\frac{\sum_{l=1}^{J}(p^{i}_{*,l}-m_{l})e_{il}}{1+\Lambda_{i}(e)}$ (A17)
$=-\frac{1}{\alpha_{i}}+R_{i}(\mathbf{p}^{i,\mathcal{M}_{k}}_{*};\mathcal{M}_{k})$
Denote $R_{i}=R_{i}(\mathbf{p}^{i,\mathcal{M}_{k}}_{*};\mathcal{M}_{k})$ as customer $i$’s optimal profit evaluated at the profit maximizing price vector of customer $i$. Multiply by $e_{ij}$ and sum over all brands $j\in[J]$:
$\sum_{j=1}^{J}(p^{i}_{*,j}-m_{j})e_{ij}$ $=-\sum_{j=1}^{J}\frac{e_{ij}}{\alpha_{i}}+\frac{\sum_{j=1}^{J}e_{ij}\left(\sum_{l=1}^{J}(p^{i}_{*,l}-m_{l})e_{il}\right)}{1+\Lambda_{i}(e)}$ (A18)
$[1+\Lambda_{i}(e)]R_{i}$ $=-\sum_{j=1}^{J}\frac{e_{ij}}{\alpha_{i}}+R_{i}\sum_{j=1}^{J}e_{ij}$
$R_{i}$ $=-\sum_{j=1}^{J}\frac{e_{ij}}{\alpha_{i}}$
$=-\sum_{j=1}^{J}\frac{\exp(\alpha_{i}p^{i}_{*,j}+v_{ij})}{\alpha_{i}}$
$=-\frac{1}{\alpha_{i}}\sum_{j=1}^{J}\exp(-1+\alpha_{i}m_{j}+v_{ij})\exp(\alpha_{i}R_{i})$
where $v_{ij}$ refers to the sum of all remaining terms that does not include price but affect the utility. Specifically, for our model $\mathcal{M}_{O}$, $v_{ij}(\mathcal{M}_{O})=\alpha_{1}\textrm{cumPercOnlineSpend}_{i}+\alpha_{3}\mathcal{I}\{\textrm{BuyLastMonth}_{ij}\}+\lambda_{1}\textrm{CF\_Price}_{j}+\lambda_{2}\textrm{CF\_Price\_cumPercOnlineSpend}_{ij}+\delta_{ij}$. For the naive model $\mathcal{M}_{N}$, $v_{ij}(\mathcal{M}_{N})=\tilde{\alpha}_{1}\mathcal{I}\{\textrm{BuyLastMonth}_{ij}\}+\tilde{\beta}_{1}\textrm{CF\_Price}_{j}+\tilde{\delta}_{ij}$. Here, for each individual customer $i$, $R_{i}=g(R_{i})$, and this equation has a unique solution since the left-hand side is increasing and starts from zero, and the right-hand side is positive at $R_{i}=0$ and decreasing. We use the binary search to obtain the optimal $R_{i}$ for each individual consumer $i$ and then calculate her profit-maximizing price vector under model $\mathcal{M}_{k}$, denoted as $\mathbf{p}^{i,\mathcal{M}_{k}}_{*}$.
F.4 Appendix to Counterfactual Scenarios
Table A21 shows the sample size for each category included in the counterfactual study (i.e., customers defined in §3.2 with a purchase occasion in June 2021).
Table A21: Sample Sizes for the Counterfactual Study - By Category
Dry Dog Food Dog Hygiene Cat Hygiene Total
Adopters 765 838 450 2425
Non-Adopters 2879 1400 1165 6314
Total 3644 2238 1615 8739

### v2

@@ A.1.2 Price and Cost Index Calculations
We construct brand-level price indices for each category by aggregating SKU-level price data from offline transactions (Gordon et al., 2013; Hitsch et al., 2021). First, we construct a store-SKU-month level price matrix from the available transactions data. To begin with, this price matrix could contain both outliers and missing values. Any observed price that is more than four times higher than or less than one-quarter of the median price for each SKU (across all store-month observations) is treated as an outlier and labeled as missing, following Hitsch et al., (2021). This process filters out unreasonable prices before imputing any missing values of store-SKU-month price (when there is no sale or there is an outlier). For these missing prices, we use the highest observed price from the previous three months for the same SKU at the same store to impute the missing value (Gordon et al., 2013). Finally, we exclude SKUs for a particular store and week if we have never observed them being sold in that store previously.
Next, we aggregate SKUs to create brand-level price indices by converting all SKU prices to a per-unit basis for comparable units. All product categories (e.g., dry food and cat sands) are measured by price per kilogram (KG) except for hygienic mats, which are measured by price per piece. Following the approach of Hitsch et al., (2021), the brand price for brand $j$ at store $s$ in month $t$, denoted as $\textrm{BrandPrice}_{jst}$, is calculated as the weighted average price per unit across all SKUs in that brand in the defined market. The weights are based on each SKU’s share of offline sales within the brand at that store over the full 30-month data period.
$\textrm{Brand\_Price}_{jst}=\sum_{k\in C_{js}}w_{jsk}\times\textrm{SKU\_Price}_{jstk},$ (A1)
where the weight of SKU $k$ is $w_{jsk}=\frac{\textrm{Store\_Sales}_{jsk}}{\sum_{k^{\prime}\in C_{js}}\textrm{Store\_Sales}_{jsk^{\prime}}}$, and $\textrm{Store\_Sales}_{jsk}$ is the total offline sales of SKU $k$ for brand $j$ at store $s$ from January 1, 2019 to June 30, 2021. $C_{js}$ represents the set of SKUs for brand $j$ available at store $s$, and $\textrm{SKU\_Price}_{jstk}$ is the price of SKU $k$ for brand $j$ sold in store $s$ during month $t$, as recorded in the store-SKU-month price matrix.
We then apply a similar approach to construct a store-SKU-month level cost matrix based on wholesale cost data, and aggregate SKU-level costs to create brand-level cost indices that correspond to each price index. These cost indices serve as the instrument for addressing price endogeneity, as discussed in Section §6. Finally, we use Brazil’s Extended National Consumer Index (IPCA) to obtain inflation-adjusted prices and costs (Brazil Institute of Geography and Statistics , 2024 IBGE), with January 2019, the first month of observation, as the base period.
@@ B.1 Supplementary Result: Percentile Price Paid
$\mathcal{I}\{\textrm{Adopter}_{i}\}\times\mathcal{I}\{t\geq T_{i}^{\textrm{adopt}}\}$ -1.437∗∗ -3.318∗∗ -0.9569 -4.104∗∗∗
@@ E.1 Detailed Results for the Role of Price Difference Across Channels
where the $\textrm{SKU\_Price}_{kstc}$ refers to the selling price of SKU $k$ of store $s$ at time $t$. The index $c$ refers to the online or offline channel. Here, a negative $\beta$ would indicate that, on average, online prices are lower than offline prices for the same SKU on the same date, and vice versa. Table A5 presents the regression result for comparing online and offline prices in Equation (A8) for three categories. Table A6 presents the regression results of decomposing SKU price variation for each category. Column (1) of Table A6 only includes the channel indicator, isOnline, and Column (2) includes both channel indicator and SKU fixed effects. Comparing the R-squared values in Table A6, we see that the price variation across channels is far smaller than the price variation across SKUs.
@@ E.2.1 Results for Alternate Demand Specification with Binary Adoption Variable
$+\alpha_{3}\mathcal{I}\{\textrm{BuyPrevOcca}_{ijt}\}+\beta_{1}\textrm{CF\_Price}_{jt}+\beta_{2}\textrm{CF\_Price\_Post\_OnlineAdoption}_{ijt}+\delta_{ij}~~\forall j\in\{1,\ldots,J\}$
To address the endogeneity of $\textrm{Price}_{jt}$ and its interaction with $\textrm{Post\_OnlineAdoption}_{it}$, we use the control function approach with instruments $\textrm{Cost}_{jt}$ and $\textrm{Cost}_{jt}\times\textrm{Post\_OnlineAdoption}_{it}$, and construct two control function terms, $\textrm{CF\_Price}_{jt}$ and $\textrm{CF\_Price\_Post\_OnlineAdoption}_{ijt}$. We then estimate the above demand model and compute the average elasticity at the customer-month level for each product category following the same steps as before. The results are shown below.
@@ E.2.2 Results for Alternate Demand Model Controlling for State Dependence due to Online Adoption
$+\alpha_{3}\mathcal{I}\{\textrm{BuyPrevOcca}_{ijt}\}+\alpha_{4}\mathcal{I}\{\textrm{BuyPrevOcca}_{ijt}\}\times\textrm{cumPercOnlineSpend}_{it}$
@@ E.3 Detailed results for testing the Parallel Pre-Trend Assumption
A fundamental condition for the credibility of the estimates presented in $\S$7.3 is the parallel trends assumption. We assess it with an event-study figure following Callaway and Sant’Anna, (2021). Figure A7 plots the point estimates and 95% bootstrapped confidence interval for each pre- and post-treatment period, and shows that all pre-treatment coefficients are insignificant from zero. We do not exhibit discernible pre-trends. In addition, the p-values from the Wald tests of the pre-trend assumption are greater than 5% level of significance in each category. We therefore conclude that the parallel trend assumption is reasonably satisfied throughout the analysis.
@@ E.4 Detailed Results for Matching, IPW Approaches to Control for Selection on Observables
[…]
@@ E.5 ATT Estimates - Using Not-Yet-Treated Adopters as Control Groups
E.6 Detailed Results for Alternative Definition of Purchase Occasions
Table A20 presents the estimates of the main demand model based on an alternative definition of the purchase occasion (i.e., customer-month level during which we observe any interaction with the retailer, either offline or online interaction). The standard errors for the demand parameters are based on the customer-level block bootstrap with 500 replications.
Table A20: Coefficients of Demand Model Estimate - Alternative Definition of Purchase Occasions
Dry Dog Food Dry Cat Food Dog Hygiene Cat Hygiene
@@ E.6 Detailed Results for Alternative Definition of Purchase Occasions
Price -0.3171∗∗∗ (0.0204) -0.1637∗∗∗ (0.0185) -1.4982∗∗∗ (0.1166) -0.2933∗∗∗ (0.0205)
cumPercOnlineSpend -1.1362∗∗ (0.365) -2.8399 (1.118) 4.822∗∗ (1.6143) 0.8407 (0.9654)
Price * cumPercOnlineSpend -0.1354∗∗∗ (0.0333) 0.0413 (0.0926) -2.8394∗∗∗ (0.7037) -0.1615 (0.0652)
BuyPrevOcca 0.173∗∗∗ (0.0263) -0.4511∗∗∗ (0.0437) -0.1633∗∗∗ (0.0275) 0.3556∗∗∗ (0.0295)
CF_Price 0.0421 (0.0257) 0.0596 (0.0247) 0.4807∗∗ (0.1714) 0.3616∗∗∗ (0.0272)
CF_Price_cumPercOnlineSpend -0.01 (0.0623) -0.1603 (0.1016) 2.372∗∗ (0.881) 0.1731 (0.0889)
Customer_Brand_FE Yes Yes Yes Yes
Observations 144,094 48,737 84,689 57,344
Log-likelihood -88735.39 -32496.48 -58287.349 -38561.677
@@
Table A21 presents the ATT estimates of offline price elasticity from CS Estimator under this alternative definition of purchase occasions and customer-month panel. The standard errors for the ATT estimates and the demand model parameters are based on the customer-level block bootstrap with 500 replications; See Figure A8.
Table A21: ATT of Online Adoption on Offline Price Elasticity (CS Estimator) – Alternative Definition of Purchase Occasions
Dependent Variable: Offline Price Elasticity
Dry Dog Food Dog Hygiene Cat Hygiene
Model: (1) (2) (3)
ATT -0.4104∗∗∗ -1.5819∗∗∗ -0.6315∗∗∗
(0.0867) (0.3688) (0.1799)
Observations 144,094 84,689 57,344
Bootstrapped standard-errors in parentheses via 500 replications
Signif. Codes: ***: 0.001, **: 0.01, *: 0.05
(a) Dry Dog Food
(b) Dog Hygiene
(c) Cat Hygiene
E.7 Detailed Results for Alternative Demand Equation with Control Function for Price Term Only
In this section, we follow Ebbes et al., (2016) to implement the control function method by running Equation (4) and adding the control function term only for price. We show that our main findings are consistent when only adding the control function terms for price. Table A22 presents the estimates of the main demand model with control function for price term only. The standard errors for the demand parameters are based on the customer-level block bootstrap with 500 replications.
Table A22: Coefficient Estimates - Demand Model with Single Control Function Term for Price
Dry Dog Food Dry Cat Food Dog Hygiene Cat Hygiene
Variables
Price -0.3165∗∗∗ (0.0212) -0.1588∗∗∗ (0.018) -1.3137∗∗∗ (0.1163) -0.2401∗∗∗ (0.0197)
cumPercOnlineSpend -0.3837 (0.4008) -1.1528 (1.1453) 3.7018∗∗ (1.1297) 2.4235 (0.9528)
Price * cumPercOnlineSpend -0.1076∗∗ (0.0374) -0.0017 (0.0974) -1.9529∗∗∗ (0.4989) -0.2061∗∗ (0.0638)
BuyPrevOcca 0.151∗∗∗ (0.0272) -0.5032∗∗∗ (0.045) -0.1805∗∗∗ (0.0283) 0.326∗∗∗ (0.0294)
CF_Price 0.0524 (0.0268) 0.0432 (0.0247) 0.3133 (0.1697) 0.3093∗∗∗ (0.0257)
Customer_Brand_FE Yes Yes Yes Yes
Observations 136,411 45,443 78,253 55,410
Log-likelihood -84211.483 -30698.359 -55455.669 -37337.88
Bootstrapped standard-errors in parentheses via 500 replications
Signif. Codes: ***: 0.001, **: 0.01, *: 0.05
Table A23 presents the ATT estimates of offline price elasticity when we estimate the demand model in Equation (7) with a control function term for price only. The standard errors for the ATT estimates and the demand model parameters are based on the customer-level block bootstrap with 500 replications; See Figure A9.
Table A23: ATT of Online Adoption on Offline Price Elasticity (CS Estimator) – Including Single CF Term for Price Endogeneity
Dependent Variable: Offline Price Elasticity
Dry Dog Food Dog Hygiene Cat Hygiene
Model: (1) (2) (3)
ATT -0.2394∗∗ -0.867∗∗∗ -0.5584∗∗∗
(0.0807) (0.2119) (0.1566)
Observations 136,411 78,253 55,410
Bootstrapped standard-errors in parentheses via 500 replications
Signif. Codes: ***: 0.001, **: 0.01, *: 0.05
(a) Dry Dog Food
(b) Dog Hygiene
(c) Cat Hygiene
[…]